# Oversmoothing, “Oversquashing”, Heterophily, Long-Range, and more: Demystifying Common Beliefs in Graph Machine Learning

### v2

@@ A.1 Deep Graph Networks
We can define a graph as a tuple $g=(\mathcal{V}_{g},\mathcal{E}_{g},\mathcal{X}_{g},\mathcal{A}_{g})$, with $\mathcal{V}_{g}$ the set of nodes, $\mathcal{E}_{g}$ the set of edges (oriented or not oriented) connecting pairs of nodes. $\mathcal{E}_{g}$ encodes the topological information of the graph and can be represented as an adjacency matrix: a binary square matrix $\mathbf{A}$ where $\mathbf{A}_{uv}$ is 1 if there is an edge between $u$ and $v$, and it is 0 otherwise. Additional node and edge features belong are represented by $\mathbf{x}_{v}\in\mathcal{X}_{g}$ and $\mathbf{a}_{uv}\in\mathcal{A}_{g}$, respectively. $\mathcal{X}_{g}$ can be, for instance, $\mathbb{R}^{d},d\in\mathbb{N}^{+}$.
@@
Most DGNs implement a synchronous message-passing mechanism, meaning each node always receives information from all neighbors at every iteration step. This local and iterative processing is at the core of DGNs’ efficiency since computation can be easily parallelized across nodes. In addition, being local means being independent of the graph’s size. When one learns the same function for all message passing iterations, we talk about recurrent architectures, as the GNN of Gori et al. [43], Scarselli et al. [85]; on the contrary, when one learns a separate parametrization for a finite number of iterations (also known as layers), we talk about convolutional architectures as the NN4G of Micheli and Sestito [71], Micheli [70].
@@
where $\mathbf{h}_{u}^{\ell}$ is the node embedding of $u$ at layer/step $\ell$, $\phi$ and $\psi$ implement learnable functions, and $\Psi$ is a permutation invariant aggregation function. Note that $\mathbf{h}^{0}_{v}=\mathbf{x}_{v}$. For instance, the Graph Convolutional Network of Kipf and Welling [54] implements the following aggregation, which is a special case of the above equation:
@@ A.2 Some OSM Definitions
Let $D=\mathrm{diag}(d_{1},\dots,d_{n})$ be the degree matrix with $d_{u}=\sum_{v}A_{uv}$. The combinatorial graph Laplacian is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, with eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$. In addition, the symmetric normalized Laplacian, defined by $\mathbf{\hat{L}}=\mathbf{I}-\mathbf{D}^{1/2}\mathbf{A}\mathbf{D}^{1/2}$, normalizes each entry to remove the influence of node degrees and its spectrum lies in $[0,2]$.
DE [26] was used to measure OSM in Cai and Wang [18] is defined as
$\hat{\mathrm{DE}}(\mathbf{H}^{\ell})=\mathrm{Tr}((\mathbf{H}^{\ell})^{T}\mathbf{\hat{L}}\mathbf{H}^{\ell})=\frac{1}{2}\sum_{u,v\in\mathcal{E}}\left\lVert\frac{\mathbf{h}_{u}}{\sqrt{d_{u}}}-\frac{\mathbf{h}_{v}}{\sqrt{d_{v}}}\right\rVert^{2}_{2}$ (4)
@@
$\mathrm{DE}(\mathbf{H}^{\ell})=\mathrm{Tr}((\mathbf{H}^{\ell})^{T}\mathbf{L}\mathbf{H}^{\ell})=\frac{1}{2}\sum_{u,v\in\mathcal{E}}\left\lVert\mathbf{h}_{u}-\mathbf{h}_{v}\right\rVert^{2}_{2}$ (5)
Other OSM metrics include with respect to the norm of node embeddings. For instance, Rayleigh Quotient (RQ) [26], which can be seen as normalized DE, is defined by
$\mathrm{RQ}=\frac{\mathrm{Tr}((\mathbf{H}^{\ell})^{T}\mathbf{\hat{L}}\mathbf{H}^{\ell})}{\lVert\mathbf{H}^{\ell}\rVert^{2}_{F}}$ (6)
first proposed for OSM analysis in [18] and later used by [104, 39, 82, 68]. RQ discerns whether embeddings become relatively smoother, independent of magnitude.
Mean Absolute Deviation (MAD) [22] averages cosine dissimilarity between a node and its neighbors.
@@
The smoothness metric SMV [59] captures a global node-distance average over all node pairs:
@@
Known theoretical bounds show $\mathrm{DE}(H^{k})$ decays exponentially with depth $k$ (e.g., the rate depends on weight spectra and graph eigenvalues) [75, 49]. Such convergence rate results primarily make assumptions on the architecture, weight matrix, and activation functions, and may be altered by skip connections, normalization layers, or simple rescaling such as $2W$ [82].
For instance Cai and Wang [18] propose a bound on the DE of two consecutive message passing layers (similar bounds found in [75, 111])
@@
being $s^{\ell}_{\max}$ the square of the maximum singular value of $W^{\ell}$, and $\lambda_{2}$ the second smallest eigenvalue of the Laplacian, i.e. the spectral gap. The proof holds when $s^{\ell}_{\max}<1/(1-\lambda_{2})$.
In addition, Giovanni et al. [39] further relate Laplacian eigenvalues and weight spectra to predict whether RQ converges to 00 (collapse) or to $\lambda_{\max}$ (no collapse).
A.3 Some OSQ Definitions
@@
A.3.1 Sensitivity
For a $K$-layer GNN let $h_{u}^{(k)}$ denote the embedding of node $u$ at layer $k$. A first proxy for OSQ is the Influence Score of Xu et al. [102],
@@ A.3.1 Sensitivity
Black et al. [16] sum these sensitivities over all unordered pairs,
@@
Giovanni et al. [39] propose a symmetric Jacobian obstruction that removes self-influence and degree bias. They define the Jacobian obstruction of node $v$ with respect to node $u$ at layer $m$ as
@@
A.3.2 Topological Bottlenecks
Many OSQ papers measure topological (structural) bottlenecks using spectral or curvature quantities. Note that here we give an intuition based on the spectral metrics [6, 52, 12], but a significant part of the literature uses metrics based on curvature [93, 60, 41, 74].
First, the topological bottleneck can be measured by Cheeger Constant [26], which is the size of the min-cut of the graph.
@@ A.3.2 Topological Bottlenecks
A small $h_{G}$ means one can separate $G$ into two large-volume parts by removing only a few edges, i.e. a severe topological bottleneck. Cheeger’s inequality links $h_{G}$ to the spectrum of G:
@@
A.3.3 Pairwise Distances
The commute time between two nodes [63] is defined as the expected number of steps that a random walker needs to go from node $u$ to $v$ and come back. The Effective Resistance between two nodes [21], $R_{uv}$, is the commute time divided by the volume of the graph [55], which is the sum of the degrees of all nodes in the graph. The effective resistance between two nodes is computed as
@@ A.3.3 Pairwise Distances
Then, some measures derived from this metric can be connected with the topological bottleneck [26, 21, 80]. For instance, the maximum effective resistance of a graph is connected with the Cheeger constant as per $R_{\max}=\max_{u,v\in\mathcal{V}}R_{uv}$
@@
In addition, the total effective resistance $R_{\text{tot}}=1/2\sum_{u,v\in\mathcal{V}}R_{uv}$ is bounded to the spectral gap [31]:
@@
A.3.4 Connecting Sensitivity and Topological Distances
The larger Total Effective Resistance ($R_{tot}=\sum_{(u,v)\in V}R_{uv}$) is, the lower the sum of pairwise sensitivities [16]:
@@ A.3.4 Connecting Sensitivity and Topological Distances
The larger the Effective Resistance is, the higher the Symmetric Jacobian Obstruction [39]:
@@ A.4 Computational Bottleneck
Following [23, 4] the receptive field was defined recursively as:
@@
which can be also seen as the set of $K$-hop neighbors neighbors, i.e. nodes that are reachable from $v$ within $K$ hops The number of nodes in each node’s receptive field can grow exponentially with the number of layers $\lvert\mathcal{N}_{v}^{K}\rvert=\mathcal{O}\left(\exp(K)\right)$ [23]. For instance, in a rooted binary tree each layer has exactly $b^{K-1}$ new neighbors, so $\lvert\mathcal{N}_{v}^{K}\rvert=1+b+b^{2}+\cdots+b^{K-1}=\Theta(b^{K})$.
When evaluating the actual computational graph resulting from message-passing, duplicates matter: a node that appears in several branches of the computation tree contributes multiple times, since each distinct walk contributes a separate message. We therefore use the multiset notation to define the computational tree for a node $v$:
@@
Therefore, we can define the notion of an "exponentially-growing receptive field" [4] as follows.
Definition A.1 (Computational Bottleneck).
@@
This definition counts every distinct $-\ell$ walk from $v$ to any node $u$. Equation (17) is exactly the row–sum of the powers of the adjacency matrix; it therefore matches the size of the computational tree.
@@
In early deep-graph networks literature, Micheli [70] introduced the idea by using the term “contextual window”: deeper layers aggregate exponentially many paths unless skip connections or global pooling curb the growth. The multiset perspective in Eq. (17) makes this explosion explicit and the matrix computation directly links to matrix-power interpretations of message passing.
Figure 6 visualises the difference between the set size $\lvert\mathcal{N}_{v}^{K}\rvert$ and the multiset size $|\mathcal{M}_{v}^{K}|$ on a toy graph and on a stochastic block model (SBM).
@@
Appendix B Sensitivity Decreases on a Grid Graph without Topological Bottlenecks
@@
Figure 7: We plot the sensitivity of the grid graph of Figure 4 for the Graph Convolutional Network [54] model for different node embedding sizes.
Appendix C Oversmoothing Does Not Always Happen
@@
Figure 8: Larger version of Figure 1. (a-b): We depict the evolution, with increasing number of layers, of the $\mathrm{DE}=\mathrm{tr}(X^{T}\Delta X)$ and the feature norm $\|X\|_{F}$, using $W$ and $2W$ feature transformations for different architectures. (c-d): Evolution of the $\mathrm{RQ}=\mathrm{tr}(X^{T}\Delta X)/\|X\|$ for $W$ and $2W$ as before. Experiments run on the Cora dataset for 50 random seeds.

### v3

@@ B.1 Deep Graph Networks
We can define a graph as a tuple $g=(\mathcal{V}_{g},\mathcal{E}_{g},\mathcal{X}_{g},\mathcal{A}_{g})$, with $\mathcal{V}_{g}$ the set of nodes, $\mathcal{E}_{g}$ the set of edges (oriented or not oriented) connecting pairs of nodes. $\mathcal{E}_{g}$ encodes the topological information of the graph and can be represented as an adjacency matrix: a binary square matrix $\mathbf{A}$ where $\mathbf{A}_{uv}$ is 1 if there is an edge between $u$ and $v$, and it is 0 otherwise. Additional node and edge features are represented by $\mathbf{x}_{v}\in\mathcal{X}_{g}$ and $\mathbf{a}_{uv}\in\mathcal{A}_{g}$, respectively. $\mathcal{X}_{g}$ can be, for instance, $\mathbb{R}^{d},d\in\mathbb{N}^{+}$.
@@
Most DGNs implement a synchronous message-passing mechanism, meaning each node always receives information from all neighbors at every iteration step. This local and iterative processing is at the core of DGNs’ efficiency since computation can be easily parallelized across nodes. In addition, being local means being independent of the graph’s size. When one learns the same function for all message passing iterations, we talk about recurrent architectures, as the GNN of Gori et al. (2005); Scarselli et al. (2009); on the contrary, when one learns a separate parametrization for a finite number of iterations (also known as layers), we talk about convolutional architectures as the NN4G of Micheli and Sestito (2005); Micheli (2009).
@@
where $\mathbf{h}_{u}^{\ell}$ is the node embedding of $u$ at layer/step $\ell$, $\phi$ and $\psi$ implement learnable functions, and $\Psi$ is a permutation invariant aggregation function. Note that $\mathbf{h}^{0}_{v}=\mathbf{x}_{v}$. For instance, the Graph Convolutional Network of Kipf and Welling (2017) implements the following aggregation, which is a special case of the above equation:
@@ B.2 Some OSM Definitions
Let $D=\mathrm{diag}(d_{1},\dots,d_{n})$ be the degree matrix with $d_{u}=\sum_{v}A_{uv}$. The combinatorial graph Laplacian is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, with eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$. In addition, the symmetric normalized Laplacian, defined by $\mathbf{\hat{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ (assuming $d_{u}>0$ for all $u$), reduces the influence of node degrees through symmetric degree normalization. For undirected graphs with nonnegative weights, the eigenvalues of $\mathbf{\hat{L}}$ lie in the interval $[0,2]$. Also, the multiplicity of eigenvalue 0 equals the number of connected components.
The Dirichlet Energy (Chung, 1997), used to measure OSM in Cai and Wang (2020), is defined as
$\hat{\mathrm{DE}}(\mathbf{H}^{\ell})=\mathrm{Tr}\left(\left(\mathbf{H}^{\ell}\right)^{T}\mathbf{\hat{L}}\mathbf{H}^{\ell}\right)=\frac{1}{2}\sum_{(u,v)\in\mathcal{E}}\left\lVert\frac{\mathbf{h}_{u}}{\sqrt{d_{u}}}-\frac{\mathbf{h}_{v}}{\sqrt{d_{v}}}\right\rVert^{2}_{2}$ (6)
@@
$\mathrm{DE}(\mathbf{H}^{\ell})=\mathrm{Tr}\left(\left(\mathbf{H}^{\ell}\right)^{T}\mathbf{L}\mathbf{H}^{\ell}\right)=\frac{1}{2}\sum_{(u,v)\in\mathcal{E}}\left\lVert\mathbf{h}_{u}-\mathbf{h}_{v}\right\rVert^{2}_{2}$ (7)
Other OSM metrics include those with respect to the norm of node embeddings. For instance, Rayleigh Quotient (RQ) (Chung, 1997), which can be interpreted as DE normalized by embedding magnitude, is defined by
$\mathrm{RQ}=\frac{\mathrm{Tr}\left(\left(\mathbf{H}^{\ell}\right)^{T}\mathbf{L}\mathbf{H}^{\ell}\right)}{\lVert\mathbf{H}^{\ell}\rVert^{2}_{F}}.$ (8)
Similarly, we can replace $\mathbf{\hat{L}}$ with the normalized Laplacian $\mathbf{L}$ to obtain the Rayleigh Quotient with respect to the normalized graph operator, denoted $\hat{\mathrm{RQ}}$. It was introduced for OSM analysis in Cai and Wang (2020) and adopted by various subsequent works (Yang et al., 2020; Di Giovanni et al., 2023b; Roth and Liebig, 2024; Maskey et al., 2023). RQ captures whether embeddings become relatively smoother, independent of magnitude.
Mean Absolute Deviation (Chen et al., 2020a) averages cosine dissimilarity between a node and its neighbors.
@@
The smoothness metric SMV (Liu et al., 2020) captures a global node-distance average over all node pairs:
@@
Known theoretical bounds show $\mathrm{DE}(H^{k})$ decays exponentially with depth $k$, where the decay rate depends on weight spectra and graph eigenvalues (Oono and Suzuki, 2020; Huang et al., 2020). Such convergence rate results primarily make assumptions on the architecture, weight matrix, and activation functions, and may be altered by skip connections, normalization layers, or simple rescaling such as $2W$ (Roth and Liebig, 2024).
For instance, Cai and Wang (2020) propose a bound on the DE of two consecutive message passing layers (similar bounds found in Oono and Suzuki (2020) and Zhou et al. (2021))
@@
being $s^{\ell}_{\max}$ the square of the maximum singular value of $W^{\ell}$, and $\lambda_{2}$ the second smallest eigenvalue of the Laplacian, i.e., the spectral gap. The proof holds when $s^{\ell}_{\max}<1/(1-\lambda_{2})$.
In addition, Di Giovanni et al. (2023b) further relate Laplacian eigenvalues and weight spectra to explain whether RQ converges to 0 (collapse) or to $\lambda_{\max}$ (no collapse).
B.3 Some OSQ Definitions
@@
B.3.1 Sensitivity
For a $K$-layer GNN let $h_{u}^{(k)}$ denote the embedding of node $u$ at layer $k$. A first proxy for OSQ is the Influence Score of Xu et al. (2018),
@@ B.3.1 Sensitivity
Black et al. (2023) sum these sensitivities over all unordered pairs,
@@
Di Giovanni et al. (2023a) propose a symmetric Jacobian obstruction that removes self-influence and degree bias. They define the Jacobian obstruction of node $v$ with respect to node $u$ at layer $m$ as
@@
B.3.2 Topological Bottlenecks
Many OSQ papers measure topological (structural) bottlenecks using spectral or curvature quantities. Note that here we give an intuition based on the spectral metrics (Arnaiz-Rodriguez et al., 2022; 2025; Karhadkar et al., 2023; Banerjee et al., 2022), but a significant part of the literature uses metrics based on curvature (Topping et al., 2022; Liu et al., 2023a; Giraldo et al., 2023; Nguyen et al., 2023).
First, the topological bottleneck can be measured by Cheeger Constant (Chung, 1997), which is the size of the min-cut of the graph.
@@ B.3.2 Topological Bottlenecks
A small $h_{G}$ means one can separate $G$ into two large-volume parts by removing only a few edges, i.e., a severe topological bottleneck. Cheeger’s inequality links $h_{G}$ to the spectrum of G:
@@
B.3.3 Pairwise Distances
The commute time between two nodes (Lovász, 1993) is defined as the expected number of steps that a random walker needs to go from node $u$ to $v$ and come back. The Effective Resistance between two nodes (Chandra et al., 1989), $R_{uv}$, is the commute time divided by the volume of the graph (Klein and Randić, 1993), which is the sum of the degrees of all nodes in the graph. The effective resistance between two nodes is computed as
@@ B.3.3 Pairwise Distances
Then, some measures derived from this metric can be connected with the topological bottleneck (Chung, 1997; Chandra et al., 1989; Qiu and Hancock, 2007). For instance, the maximum effective resistance of a graph is connected with the Cheeger constant as per $R_{\max}=\max_{u,v\in\mathcal{V}}R_{uv}$
@@
In addition, the total effective resistance $R_{\text{tot}}=1/2\sum_{u,v\in\mathcal{V}}R_{uv}$ is bounded to the spectral gap (Ellens et al., 2011):
@@
B.3.4 Connecting Sensitivity and Topological Distances
The larger Total Effective Resistance ($R_{tot}=\sum_{(u,v)\in V}R_{uv}$) is, the lower the sum of pairwise sensitivities (Black et al., 2023):
@@ B.3.4 Connecting Sensitivity and Topological Distances
The larger the Effective Resistance is, the higher the Symmetric Jacobian Obstruction (Di Giovanni et al., 2023a):
@@ B.4 Computational Bottleneck
Following Chen et al. (2018) and Alon and Yahav (2021) the receptive field was defined recursively as:
@@
which can be also seen as the set of $K$-hop neighbors, i.e., nodes that are reachable from $v$ within $K$ hops The number of nodes in each node’s receptive field can grow exponentially with the number of layers $\lvert\mathcal{N}_{v}^{K}\rvert=\mathcal{O}\left(\exp(K)\right)$ (Chen et al., 2018). For instance, in a rooted binary tree each layer has exactly $b^{K-1}$ new neighbors, so $\lvert\mathcal{N}_{v}^{K}\rvert=1+b+b^{2}+\cdots+b^{K-1}=\Theta(b^{K})$.
Computational Tree
When evaluating the actual computational graph resulting from message-passing, duplicates matter: a node that appears in several branches of the computation tree contributes multiple times, since each distinct walk contributes a separate message. We therefore use the multiset notation to define the computational tree for a node $v$ in Eq. 3, reproduced here for convenience:
@@
Computational Bottleneck
Therefore, we defined in Def. 4.1 the computational bottleneck of node $v$ as the size of the computational tree, $|\mathcal{M}^{K}_{v}|$, for a given node $v$ and number of message passing layers $K$.
@@
This definition counts every distinct length-$\ell$ walk from $v$ to any node $u$. Equation equation 19 is exactly the row–sum of the powers of the adjacency matrix; it therefore matches the size of the computational tree.
@@
In early deep-graph networks literature, Micheli (2009) introduced the idea by using the term “contextual window”: deeper layers aggregate exponentially many paths unless skip connections or global pooling curb the growth. The multiset perspective in Eq. equation 19 makes this explosion explicit and the matrix computation directly links to matrix-power interpretations of message passing.
Figure 6 visualises the difference between the set size $\lvert\mathcal{N}_{v}^{K}\rvert$ and the multiset size $|\mathcal{M}_{v}^{K}|$ on a toy graph.
@@
Appendix C Sensitivity Decreases on a Grid Graph without Topological Bottlenecks
@@
Figure 7: We plot the sensitivity of the grid graph of Figure 4 for the Graph Convolutional Network (Kipf and Welling, 2017) model for different node embedding sizes.
Appendix D Oversmoothing Does Not Always Happen
@@
Figure 8: Larger version of Figure 1. (a-b): We depict the evolution, with increasing number of layers, of the $\mathrm{DE}$ and the feature norm $\|X\|_{F}$, using $W$ and $2W$ feature transformations for different architectures. (c-d): Evolution of the $\mathrm{RQ}$ for $W$ and $2W$ as before. Experiments run on the Cora dataset for 50 random seeds.
Appendix E Alternative Views
One may argue that the common beliefs of Table 1 could be regarded as alternative views, which we have abundantly discussed and challenged in previous sections. However, we additionally discuss other alternative viewpoints.
One is put forward by Blayney et al. (2025), which states that the oversquashing term is still useful to describe issues arising from topological bottlenecks and the depth in the computational tree. The authors argue that all performance issues related to “oversquashing” are caused by either model capacity or low sensitivity, and we do not intend to contest their observations in this space. Their alternative position on the use of the “oversquashing” term likely stems from an in-depth understanding of the distinction between topological and computational bottlenecks as well as the underlying assumptions, which, as this paper shows, cannot be applied to the whole community yet. As a matter of fact, our proposal is pedagogical rather than practical: Section 4 does not propose how to identify specific problems (e.g, capacity or sensitivity), but how to better think about and approach them without confusing everything under a single umbrella term. For this reason, we believe that a clear separation of concepts, rather than the use of the term “oversquashing”, can have a profound impact on the whole community in formulating more precise research questions.
Another alternative view is the recent one of Mishayev et al. (2025), which proposes, similarly to Blayney et al. (2025) a problem-oriented dichotomy for the “oversquashing” phenomenon. They argue for the decoupling of this term into computational bottlenecks, which can happen in short-range tasks, and vanishing gradients, which seem to be more related to long-range dependencies. However, because computational bottlenecks can appear in long-range tasks as well as vanishing gradients, we believe this work focuses more on identifying practical problems to address rather than providing a conceptual separation of concepts related to oversquashing. One evidence for this is that the synthetic tasks used in the paper, namely two-radius and ring transfer, contain a topological bottleneck that may have been important to reaching some conclusions. Hence, our goal is complementary to that of Mishayev et al. (2025).
Finally, we mention the position of Kormann et al. (2026). The authors highlight that improvements in over-smoothing metrics and model performance are, at best, weakly correlated. The hypothesis they investigate is that most real tasks rely on short-range dependencies. In addition, they question the focus on curvature in OSQ contributions and call for a more principled understanding supported by appropriate statistics. The authors’ views and empirical observations are therefore complementary to ours.